OpenStudy (yuki):

I will give you 20 medals if you could help me solve this linear algebra problem.

OpenStudy (kira_yamato):

What is it?

OpenStudy (yuki):

Thank god finally

OpenStudy (yuki):

\[T(\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]) = \left[\begin{matrix}10a-9b & 4a-2b \\ -2c+5d & c+2d\end{matrix}\right]\]

OpenStudy (yuki):

I need to find the matrix that represents T

OpenStudy (kira_yamato):

Check working before use...

OpenStudy (yuki):

your multiplication is not on the correct side

OpenStudy (yuki):

(a b : c d) is multiplied on the right, so it's inverse is also multiplied on the right

OpenStudy (yuki):

Man, the only day where no one likes to do linear algebra

OpenStudy (yuki):

Kira, although that is not the answer I am looking for you deserve to have the medals because you helped. Thank you

OpenStudy (kira_yamato):

ありがとう

OpenStudy (yuki):

いいえこちらこそ